# Disentangling Boltzmann brains, the time-asymmetry of memory, and the second law

### v3

@@ IV Discussion
We will emphasize that since the entropy conjecture cannot make a conclusion regarding probability of the BB hypothesis that doesn’t resort to priors (see Appendix), any conclusion that the BB hypothesis should be discarded entirely must move to the level of theory building and/or justification of priors. And as stated previously, it seems very hard to justify a prior that allows one to conclude that the BB hypothesis is true. In the context of theory building, philosophy of science has provided no shortage of qualities that might help define a “good” theory. Cognitive stability Carroll (2020) is one such quality. Yet, critically, there is not a fully rigorous argument, yet, in the domain of physics to dispel the possibility of the BB hypothesis. The view we have established in this paper is that the BB hypothesis is not established by the standard argument, nor is it refuted by standard criticisms. It seems the BB hypothesis falls out of the domain of science, into the grey ocean of hypothesis described in the introduction. We cannot rigorously assign probability to something unknowable, nor can we say that something unknowable is not true.
It is important to appreciate that there are many deep issues in the philosophy of science that our careful investigation of the BB hypothesis uncovers. One is that there is the huge issue of how — from the perspectives of either philosophy or the axiomatizations of Bayesian probability — one can justify conditioning on any moment in addition to the present, and if so, why just one such additional moment.
@@ B.2 Unstable reasoning in analysis of Boltzmann Brains
Stated more prosaically, yes, under the BB hypothesis the present states of our memory systems might just be neurobiological functions, completely unreliable in that they don’t imply what they seem to. But it’s also quite possible under the BB hypothesis that those memory systems of our past are highly reliable, even though (under the BB hypothesis) the entropy of the universe as a whole is far higher in the past than the present.

### v4

@@ V Discussion
We emphasize that since the entropy conjecture cannot make a conclusion regarding the probability of the BB hypothesis that doesn’t resort to priors (see Appendix), any conclusion that the BB hypothesis should be discarded entirely must involve justification of priors. And as stated previously, it seems very hard to justify a prior that allows one to conclude that the BB hypothesis is not true. There is not a fully rigorous argument, yet, in the domain of physics to dispel the possibility of the BB hypothesis. The view we have established in this paper is that the BB hypothesis is not established by the standard argument, nor is it refuted by standard criticisms.
It is important to appreciate that there are many deep issues in the philosophy of science that our careful investigation of the BB hypothesis uncovers. One is the huge issue of how — from the perspectives of either philosophy or the axiomatizations of Bayesian probability — one can justify conditioning on any moment in addition to the present, and if so, why just one such additional moment.
@@ B.2 Unstable reasoning in analysis of Boltzmann Brains
Stated more prosaically, yes, under the BB hypothesis the present states of our memory systems might just be neurobiological fluctuations, completely unreliable in that they don’t imply what they seem to. But it’s also quite possible under the BB hypothesis that those memory systems of our past are highly reliable, even though (under the BB hypothesis) the entropy of the universe as a whole is far higher in the past than the present.